# Accurate stochastic simulation algorithm for multiscale models of infectious diseases

### v1

@@ 2.1.4 Model coupling
With the specific population-level and within-host sub-models discussed in Subsection 2.1.1 and Subsection 2.1.3, respectively, we can now construct a multiscale model by linking these separate sub-systems. Theoretically, this coupling process can be facilitated through identifying possible parameters and variables in one sub-system that affect the dynamics of the other and consistently formulating a feedback across sub-models. A common approach to accomplish this is by expressing that set of parameters or variables in one sub-model as functions of those in the other [GMN14]. In this Subsection, we will develop a test multiscale model that links the population-level transmission coefficient to the within-host viral load. It is important to note that, to illustrate the simulation methodology, the model coupling presented below represents just one method to connect these scales.
@@
where $v(\delta t)=lV(\delta t)$. In (15), $t_{j}$ is the calendar time associated with the initial infection of individual $j$ and $\delta t_{j}$ is the age of the infection in that individual. Note that the individuality of $\beta_{j}$ is encapsulated only in the initial infection time $t_{j}$. We further assume that within-host dynamics is independent from the population-level dynamics (and can be solved numerically once at the start of the simulation). The function $v(\delta t)$ is assumed to be zero prior to infection ($\delta t<0$), so as to account for the agent being susceptible (noninfectious).
@@
Since the propensity $\alpha^{(i)}$ now depends explicitly on time, the previously mentioned algorithms in Subsection 2.1.2 need to be updated. Furthermore, it is important to note that this function of time is complex as the infection times $\{t_{j}\}$ are stochastic.
@@
Figure 2: (a) Schematic diagram for the within-host model. The boxes represent different cell compartments while the arrows denote transitions between compartments. Note that the compartments linked to the tail end of the arrow, ()maps-to($\mapsto$)( ↦ ), does not experience any cell nor virus flow in the corresponding transition. The dashed arrows denote the removal rates. (b) Sample numerical solution of the within-host model (14). The solution is plotted as a function of infection age $\delta t$ from MATLAB’s $ode15s$ solver with both absolute and relative tolerances being $10^{-8}$. We set the initial condition $\left(C_{0},C_{0}^{*},V_{0}\right)$ as $(10^{8},10^{0},10^{6})$ number of cells or viral particles. The parameter values chosen are: $p=10\rm\text{ }day^{-1}$, $c=1\rm\text{ }day^{-1}$, $k=10^{-7}\rm\text{ }day^{-1}\rm cell^{-1}$, $\mu_{c}=10^{-1}\rm\text{ }day^{-1}$, $\delta_{c}=10\rm\text{ }day^{-1}$, and $\Lambda_{c}=(1.111\times 10^{6})\text{ cell}\times\text{day}^{-1}$. (c) Viral dynamics for each of 5 individuals. These plots also represent the per capita forces of infection ($v(t-t_{j})$) divided by the coupling constant ($l$) for each of 5 individuals (see (15)), with different calendar infection dates $\{t_{j}\}=\{19,78,18,82,43\}$ respectively generated by the population-level SSA. They are shown relative to the calendar time $t$. (d) Per susceptible propensity of infection. The propensity of an infection per infected individual $\alpha^{(i)}(t)/s$ in the population-level model assumes the complex form of the sum of each individual transmission coefficient by (16). We choose the coupling constant $l$ to take the value of one per day in this plot.
@@ 2.2 Multiscale model simulation
The within-host model (14) can be solved numerically to give $v(\delta t)$. The multiscale stochastic model can therefore be written in the same manner as (3) in Subsection 2.1.2, but noting the change in transmission coefficient as shown in (16). Probabilities for updating the simulation from known state $(s,i)$ from (4)-(8) in the multiscale case at time $t$ is conditional on the set of times in the past individuals firstly became infectious. That is, by denoting $P_{(s,i)}(t)$ as the probability of being in state $(s,i)$ conditional on known infection times $\{t_{j}\}_{j=1}^{i}$:
@@
$P_{(s-1,i+1)}(t+\Delta t)$ $=\alpha^{(i)}(s,i;t)\Delta t:=s\sum_{j=1}^{i}v(t-t_{j})\Delta t.$ (21)
Applying the NRM in the case of the time dependent total propensity $\alpha(s,i;t)$, putative times for birth and deaths can still be found in the standard way as their Markovian property still holds. However, drawing a putative time for the next infection event requires sampling from a distribution with a time dependent rate. The derivation of such putative time is to sample the time for a particular event conditional on the state remaining unchanged. According to (21), the probability of an infection in time interval $(t,t+\Delta t)$ is $s\Delta t\sum_{j=1}^{i}v(t-t_{j})+O(\Delta t^{2})$. Let $\tilde{p}(n)$ be the probability that no infection occurs in the first $n$ time steps from $t$. We have
@@
A sample of the putative time for an infection $\tau^{(i)}$ can be generated by sampling a random number $u\sim U(0,1)$ and solving $\text{CDF}_{s}\left(\tau^{(i)};t\right)=u$ for $\tau^{(i)}$, where $t$ is the current calendar time. This is complicated because the exponent in $\text{CDF}_{s}\left(\tau^{(i)};t\right)$ as shown in (23) is a nontrivial function which depends on the stochastic history of infections $\{t_{j}\}_{j=1}^{i}$ among the infectious population. Furthermore, even if efficient sampling from the CDF in (23) is solved, $\tau^{(i)}$ is only a putative time for infection at current calendar time $t$; once the within-host as well as the population-level states change when time forwards and when any event occurs, a new putative time would need to be sampled. In Subsection 2.2.1, we will firstly outline an approximate computational algorithm that uses discrete time steps $\Delta t$ and equations (17)-(21) to simulate the multiscale scenario. Since the algorithm propagates forward in time using fixed time steps, they are usually called ‘time-driven’ algorithms. This is in contrast to the direct SSA or NRM approaches which propagate forward in time by the occurrence of events; aptly named ‘event-driven’ approaches. We will then outline our main result in Subsection 2.2.2; an exact NRM method where $\tau^{(i)}$ is efficiently sampled using the CDF presented in (23).
@@ 2.2.1 Approximate time-driven simulation
The time-driven algorithm we will present in this subsection is rather simple but is only approximate. It uses a predetermined fixed small time step $\Delta t$ with brute force to simulate events using (17)-(21). Reducing the size of $\Delta t$ allows for control over the level of accuracy but at the cost of efficiency. We shall use this algorithm to compare with the main algorithm of this paper in Subsection 2.2.2. The pseudocode for this algorithm is shown in Algorithm 1. The algorithm is a simple extension of the time-driven algorithm previously discussed in Subsection 2.1.2.
@@
8:Initialise the fixed time stepping size $\Delta t$ (for accurate results ensure that probabilities in (18)-(21) are always small (less than some tolerance $\epsilon$): $\max\left\{\alpha_{b}+\alpha_{ds}+\alpha_{di}+s\sum_{j=1}^{i}v(t-t_{j})\right\}\Delta t\leq\epsilon,\,\forall t$, where $v(\delta t)$ is defined in (15)).
@@ 2.2.2 Accurate multiscale simulation
In NRM, the event with the minimum putative time is selected and the state $(s,i)$ is updated to $(\bar{s},\bar{i})$ corresponding to the change caused by that event. After changing the state, the time is first updated and then the putative time for the event that occurred is resampled. For all other events, the putative time is updated to reflect the progression of time and then rescaled to reflect any change in propensity as a result of the transition $(s,i)$ to $(\bar{s},\bar{i})$. In this section, we present a NRM algorithm for simulating non-Markovian disease transmission models of the type explored in this manuscript. We focus on how putative times corresponding to within-host infections $\tau^{(i)}$ are computed, resampled, and rescaled to be used in the general next reaction method framework.
To find $\tau^{(i)}$ we could attempt to solve $\mathrm{SDF}_{s}(\tau^{(i)};t)=u$ where $u=1-\tilde{u}\sim U(0,1)$, where $\mathrm{SDF}_{s}$ is defined in (23). However, instead of finding a putative time $\tau^{(i)}$ for the next infection, we note that infected individuals are unique in this model (at least when it comes to infecting others) and therefore it is more tractable to subdivide infection events by the individual who is responsible for the infection; and find a putative infection time for each infected individual $j$ as $\tau_{j}^{(i)}$. We shall then use the minimum of these putative times by the NRM to determine $\tau^{(i)}=\mathrm{min}_{j}(\tau_{j}^{(i)})$. In fact, going further than this we determine the update time $\tau^{(m)}=\mathrm{min}(\{\tau_{j}^{(i)}\},\tau^{(di)},\tau^{(ds)},\tau^{(b)})$. Using this method of treating each individual infected person separately, we also have the capability of storing data on who is responsible for distributing the disease for later analysis without any further calculation.
We are permitted to (and indeed benefited by) applying the NRM in this way because
@@
and similarly $\text{CDF}_{s;j}=1-\text{SDF}_{s;j}$. Note that $1\leq j\leq i$; $t-t_{j}$ denotes the age of infection; $\Psi$ is defined in (15). Let $\tau_{m}^{(i)}=\mathrm{min}_{j}(\tau_{j}^{(i)})$. The NRM asserts that $\tau_{m}^{(i)}$ and $\tau^{(i)}$ have the same distribution, which can easily be seen by (23)
@@
In the NRM, when $\tau^{(m)}=\tau_{j}^{(i)}$, the ‘next event’ is infection by the $j$th infected individual. For this individual, the next infection will need to be resampled. For all other putative times, $\tau_{k}^{(i)}<\tau^{(m)}$ where $k\neq j$, the NRM describes a rescaling of the putative time instead of resampling for efficiency. We describe both resamplng and rescaling next in detail.
@@
That is, the value of $\psi_{j}=-\log(u_{j})/s$ gives the putative increase in the function $\Psi$ (which is independent of $s$) from the current time to the time of the infection. We therefore track four values for each infected individual: (1) the current value of the individuals infection time $\delta t_{j}$, (2) the current putative time until the individual next infects a susceptible $\tau_{j}^{(i)}$, (3) the current value of $\Psi_{j}=\Psi(\delta_{j})$, and (4) the current putative change the value of $\Psi_{j}$ when the individual next infects $\psi_{j}=\Psi(\tau_{j}^{(i)}+\delta t_{j})-\Psi_{j}$. It is the later putative variable $\psi_{j}=-\log(u_{j})/s$ (not the putative time $\tau_{j}^{(i)}$) which can easily be sampled. However, as time moves forward and infection time needs updating, this has to be done in the time domain by shifting $\delta t_{j}$. Resampling should be done if $\tau^{(m)}=\tau_{j}^{(i)}$.
@@
where the bar notation indicates the ‘new’ value of the variable. The update step also allows for the calculation of the new $\Psi_{j}$, that is $\bar{\Psi_{j}}=\Psi(\bar{\delta t_{j}})$ by the lookup table for the function $\Psi$.
Second, after the change of state $(s,i)$ to $(\bar{s},\bar{i})$ a new putative $\psi$ is sampled for the individual; $\bar{\psi_{j}}=-\log{u_{j}}/\bar{s}$ by (27) reflecting the new $\bar{s}$. Finally, it is possible to determine the new putative time $\bar{\tau_{j}^{(i)}}$. This is done simply by knowing $\bar{\delta t_{j}}$, $\bar{\Psi_{j}}$ and $\bar{\psi_{j}}$ and using the lookup table to invert the function $\Psi$:
@@
A schematic showing how $\tau^{(i)}_{j}$ is resampled is presented in Figure 3. Here we show how the update in time (green) is to the putative time for the individual $j$ ($\tau^{(m)}=\tau^{(i)}_{j}$) and the time is first updated. By the green dashed line, it can be seen how the new infection time for the $j$th individual, $\bar{\delta t}_{j}$, corresponds to the new value $\bar{\Psi}_{j}$ for this infectious individual. From here, the resample step is given in red. A putative future increment in $\Psi$ ($\psi_{j}=-\log(u_{j})/s$) is computed and then (by the red dashed line) a new putative time until next infection $\bar{\tau}_{j}^{(i)}$ is determined.
@@
If $\tau^{(m)}<\tau_{j}^{(i)}$ naively, it is possible to simply update $\delta t_{j}$, $\tau_{j}^{(i)}$, $\Psi_{j}$ and $\psi_{j}$ using the resampling method. This is however costly because it involves generating a new random number $u_{j}$. Instead, a rescaling is possible. But unlike the classical NRM where this rescaling is done to $\tau_{j}^{(i)}$, it is necessary to instead rescale in $\psi_{j}$. To be clear, $\bar{\delta t_{j}}$ and $\bar{\Psi_{j}}$ are calculated as before. We note however that since $\tau_{j}^{(i)}>\tau^{(m)}$, there is a portion of time between $\bar{\delta t_{j}}$ and $\delta t_{j}+\tau_{j}^{(i)}$ for which the propensity is changed due to the change of state $\bar{s}$ from its original value $s$. This corresponds to the interval in $\Psi$ of width

### v2

@@ 2.1.4 Model coupling
With the specific population-level and within-host sub-models discussed in Section 2.1.1 and Section 2.1.3, respectively, we can now construct a multiscale model by linking these separate sub-systems. Theoretically, this coupling process can be facilitated through identifying possible parameters and variables in one sub-system that affect the dynamics of the other and consistently formulating a feedback across sub-models. A common approach to accomplish this is by expressing that set of parameters or variables in one sub-model as functions of those in the other [GMN14]. In this section, we will develop a test multiscale model that links the population-level transmission coefficient to the within-host viral load. It is important to note that, to illustrate the simulation methodology, the model coupling presented below represents just one method to connect these scales.
@@
where $v(t-t_{j})=lV(\delta t_{j})$ and $t-t_{j}=\delta t_{j}$. In (14), $t_{j}$ is the calendar time associated with the initial infection of individual $j$ and $\delta t_{j}$ is the age of the infection in that individual. Note that the individuality of $\beta_{j}$ is encapsulated only in the initial infection time $t_{j}$. We further assume that within-host dynamics are independent of the population-level dynamics (and can be solved numerically once at the start of the simulation). The function $v(\delta t)$ is assumed to be zero prior to infection ($\delta t<0$), so as to account for the agent being susceptible (noninfectious).
@@
Since the propensity $\alpha^{(i)}$ now depends explicitly on time, the previously mentioned algorithm in Section 2.1.2 needs to be updated. Furthermore, it is important to note that this function does not have a simple closed-form expression, as the infection times $\{t_{j}\}$ are stochastic.
@@
Figure 2: (a) Schematic diagram for the within-host model. The boxes represent different cell compartments while the arrows denote transitions between compartments. Note that the compartments linked to the tail end of the arrow, ()maps-to($\mapsto$)( ↦ ), does not experience any cell nor virus flow in the corresponding transition. The dashed arrows denote the removal rates. (b) Sample numerical solution of the within-host model (13). The solution is plotted as a function of infection age $\delta t$ from MATLAB’s $ode15s$ solver with both absolute and relative tolerances being $10^{-8}$. We set the initial condition $\left(C_{0},C_{0}^{*},V_{0}\right)$ as $(10^{8},10^{0},10^{6})$ number of cells or viral particles. The parameter values chosen are: $p=10\rm\text{ }day^{-1}$, $c=1\rm\text{ }day^{-1}$, $k=10^{-7}\rm\text{ }day^{-1}\rm cell^{-1}$, $\mu_{c}=10^{-1}\rm\text{ }day^{-1}$, $\delta_{c}=10\rm\text{ }day^{-1}$, and $\Lambda_{c}=(1.111\times 10^{6})\text{ cell}\times\text{day}^{-1}$(c) Viral dynamics for each of 5 individuals. These plots also represent the per capita forces of infection ($v(t-t_{j})$) divided by the coupling constant ($l$) for each of 5 individuals (see (14)), with different calendar infection dates $\{t_{j}\}=\{19,78,18,82,43\}$ respectively generated by the population-level SSA. They are shown relative to the calendar time $t$(d) Per susceptible propensity of infection. The propensity of an infection per infected individual $\alpha^{(i)}(t)/s$ in the population-level model assumes the complex form of the sum of each individual transmission coefficient by (15). We choose the coupling constant $l$ to take the value of one per day in this plot.
@@ 2.2 Multiscale model simulation
The within-host model (13) can be solved numerically to give $v(\delta t)$. The multiscale stochastic model can therefore be written in the same manner as (3) in Section 2.1.2, but noting the change in transmission coefficient as shown in (15). Probabilities for updating the simulation from known state $(s,i)$ from (4)-(8) in the multiscale case at time $t$ is conditional on the set of times in the past individuals firstly became infectious. That is, by denoting $P_{(s,i)}(t)$ as the probability of being in state $(s,i)$ conditional on known infection times $\{t_{j}\}_{j=1}^{i}$:
@@
$P_{(s-1,i+1)}(t+\Delta t)$ $=\alpha^{(i)}(s,i;t)\Delta t=s\sum_{j=1}^{i}v(t-t_{j})\Delta t.$ (20)
Applying the strategy used in NRM in the case of the time dependent total propensity $\alpha(s,i;t)$, putative times for birth and deaths can still be found in the standard way as their Markovian property still holds. However, drawing a putative time for the next infection event requires sampling from a distribution with a time dependent rate. The derivation of such putative time is to sample the time for a particular event conditional on the state remaining unchanged. According to (20), the probability of an infection in time interval $(t,t+\Delta t)$ is $s\Delta t\sum_{j=1}^{i}v(t-t_{j})+O(\Delta t^{2})$. Let $\tilde{p}(n)$ be the probability that no infection occurs in the first $n$ time steps from $t$. We have
@@
A sample of the putative time for an infection $\tau^{(i)}$ can be generated by sampling a random number $u\sim U(0,1)$ and solving $\text{CDF}_{s}\left(\tau^{(i)};t\right)=u$ for $\tau^{(i)}$, where $t$ is the current calendar time. This is complicated because the exponent in $\text{CDF}_{s}\left(\tau^{(i)};t\right)$ as shown in (22) is a nontrivial function which depends on the stochastic history of infections $\{t_{j}\}_{j=1}^{i}$ among the infectious population. Furthermore, even if efficient sampling from the CDF in (22) is solved, $\tau^{(i)}$ is only a putative time for infection at current calendar time $t$; once the within-host as well as the population-level states change when time forwards and when any event occurs, a new putative time would need to be sampled. In Section 2.2.1, we will firstly outline an approximate computational algorithm that uses discrete time steps $\Delta t$ and equations (16)-(20) to simulate the multiscale scenario. Since the algorithm propagates forward in time using fixed time steps, they are usually called ‘time-driven’ algorithms. This is in contrast to the NRM approach which propagates forward in time by the occurrence of events; aptly named ‘event-driven’ approaches. We will then outline our main result in Section 2.2.2, which uses a similar strategy as in NRM method where $\tau^{(i)}$ is appropriately resampled and rescaled using the CDF presented in (22).
@@ 2.2.1 Approximate time-driven simulation
The time-driven algorithm we will present in this section is rather simple but is only approximate. It uses a predetermined fixed small time step $\Delta t$ with brute force to simulate events using (16)-(20). Reducing the size of $\Delta t$ allows for control over the level of accuracy but at the cost of efficiency. The algorithm is a simple extension of the time-driven algorithm previously discussed in Section 2.1.2. We shall use this algorithm to compare with the main algorithm of this paper in Section 2.2.2. The pseudocode for this algorithm is shown in Algorithm 1, whereas the MATLAB implementation and all data generated from the simulations are available on Github at https://github.com/YuanYIN99/Accurate-stochastic-simulation-algorithm-for-multiscale-models-of-infectious-diseases.git.
@@
8:Initialise the fixed time stepping size $\Delta t$ (for accurate results ensure that probabilities in (17)-(20) are always small (less than some tolerance $\epsilon$): $\max\left\{\alpha^{(b)}+\alpha^{(ds)}+\alpha^{(di)}+s\sum_{j=1}^{i}v(t-t_{j})\right\}\Delta t\leq\epsilon,\,\forall t$, where $v(\delta t)$ is defined in (14)).
@@ 2.2.2 Accurate multiscale simulation
Based on the strategy in NRM, the event with the minimum putative time is selected and the state $(s,i)$ is updated to $(\bar{s},\bar{i})$ corresponding to the change caused by that event. After changing the state, the time is first updated and then the putative time for the event that occurred is resampled. For all other events, the putative time is updated to reflect the time that has passed and then rescaled to reflect any change in propensity as a result of the transition $(s,i)$ to $(\bar{s},\bar{i})$. In this section, we present an algorithm based on the idea of NRM, for simulating non-Markovian disease transmission models of the type explored in this manuscript. We focus on how putative times corresponding to within-host infections $\tau^{(i)}$ are computed, resampled, and rescaled to be used in the general next reaction method framework.
To find $\tau^{(i)}$ we could attempt to solve $\mathrm{SDF}_{s}(\tau^{(i)};t)=u$ where $u=1-\tilde{u}\sim U(0,1)$, where $\mathrm{SDF}_{s}$ is defined in (22). However, instead of finding a putative time $\tau^{(i)}$ for the next infection, we note that infected individuals are unique in this model (at least when it comes to infecting others) and therefore it is more tractable to subdivide infection events by the individual who is responsible for the infection; and find a putative infection time for each infected individual $j$ as $\tau_{j}^{(i)}$. We shall then use the minimum of these putative times, based on the idea from NRM, to determine $\tau^{(i)}=\mathrm{min}_{j}(\tau_{j}^{(i)})$. In fact, going further than this we determine the update time $\tau^{(m)}=\mathrm{min}(\{\tau_{j}^{(i)}\},\tau^{(di)},\tau^{(ds)},\tau^{(b)})$. Using this method of treating each individual infected person separately, we also have the capability of storing data on who is responsible for distributing the disease for later analysis without any further calculation.
We are permitted to (and indeed benefited by) applying the strategy of NRM in this way because
@@
and similarly $\text{CDF}_{s;j}=1-\text{SDF}_{s;j}$. Note that $1\leq j\leq i$; $t-t_{j}$ denotes the age of infection; $\Psi$ is defined in (23). Let $\tau_{m}^{(i)}=\mathrm{min}_{j}(\tau_{j}^{(i)})$. The idea of NRM asserts that $\tau_{m}^{(i)}$ and $\tau^{(i)}$ have the same distribution, which can easily be seen by
@@
As in the NRM, when $\tau^{(m)}=\tau_{j}^{(i)}$, the ‘next event’ is infection by the $j$th infected individual. For this individual, the next infection will need to be resampled. For all the other putative times, $\tau_{k}^{(i)}>\tau^{(m)}$ where $k\neq j$, we apply a similar strategy as in NRM for rescaling of the putative time instead of resampling. We describe both resampling and rescaling next in detail.
@@
That is, the value of $\psi_{j}=-\log(u_{j})/s$ gives the putative increase in the function $\Psi$ (which is independent of $s$) from the current time to the time of the infection. We therefore track four values for each infected individual: (1) the current value of the individuals infection time $\delta t_{j}$, (2) the current putative time until the individual next infects a susceptible $\tau_{j}^{(i)}$, (3) the current value of $\Psi_{j}=\Psi(\delta t_{j})$, and (4) the current putative change the value of $\Psi_{j}$ when the individual next infects $\psi_{j}=\Psi(\tau_{j}^{(i)}+\delta t_{j})-\Psi_{j}(\delta t_{j})$. It is the later putative variable $\psi_{j}=-\log(u_{j})/s$ (not the putative time $\tau_{j}^{(i)}$) which can easily be sampled. However, as time moves forward and infection time needs updating, this has to be done in the time domain by shifting $\delta t_{j}$. Resampling should be done if $\tau^{(m)}=\tau_{j}^{(i)}$.
@@
where the bar notation indicates the ‘new’ value of the variable. The update step also allows for the calculation of the new $\Psi_{j}$, that is $\bar{\Psi_{j}}=\Psi(\bar{\delta t_{j}})$ by the look-up table for the function $\Psi$. Second, after the change of state $(s,i)$ to $(\bar{s},\bar{i})$ a new putative $\psi$ is sampled for the individual; $\bar{\psi_{j}}=-\log{u_{j}}/\bar{s}$ by (26) reflecting the new $\bar{s}$. Finally, it is possible to determine the new putative time $\bar{\tau_{j}^{(i)}}$. This is done simply by knowing $\bar{\delta t_{j}}$, $\bar{\Psi_{j}}$ and $\bar{\psi_{j}}$ and using the look-up table to invert the function $\Psi$:
@@
A schematic showing how $\tau^{(i)}_{j}$ is resampled is presented in Figure 3. Here we show how the update in time (green) is to the putative time for the individual $j$ ($\tau^{(m)}=\tau^{(i)}_{j}$) and the time is first updated. By the green dashed line, it can be seen how the new infection time for the $j$th individual, $\bar{\delta t}_{j}$, corresponds to the new value $\bar{\Psi}_{j}$ for this infectious individual. From here, the resample step is given in red. A putative future increment in $\Psi$ ($\bar{\psi_{j}}=-\log(u_{j})/\bar{s}$) is computed and then (by the red dashed line) a new putative time until next infection $\bar{\tau}_{j}^{(i)}$ is determined.
@@
If $\tau^{(m)}<\tau_{j}^{(i)}$ naively, it is possible to simply update $\delta t_{j}$, $\tau_{j}^{(i)}$, $\Psi_{j}$ and $\psi_{j}$ using the resampling method. Whilst random number generation is cheap in comparison to looking up the value of a nonlinear function and its inverse, it is possible to determine the new putative time $\tau_{j}^{(i)}$ without generating any new random numbers and only at the cost of a few extra floating point operations. This can be achieved by simply rescaling $\psi_{j}$. We begin by determining, $\bar{\delta t_{j}}$ and $\bar{\Psi_{j}}$ as previously described. We note however that since $\tau_{j}^{(i)}>\tau^{(m)}$, there is a portion of time between $\bar{\delta t_{j}}$ and $\delta t_{j}+\tau_{j}^{(i)}$ for which the propensity is changed due to the change of state $\bar{s}$ from its original value $s$. This corresponds to the interval in $\Psi$ of width